# RAD: Training an End-to-End Driving Policy via Large-Scale 3DGS-based Reinforcement Learning

### v1

@@ 3.5 Policy Optimization
0:1 0.229 0.211 0.018 0.066 0.039 0.027 0.238 3.928 0.103
1:0 0.143 0.128 0.015 0.080 0.065 0.015 0.345 4.204 0.085
2:1 0.137 0.125 0.012 0.059 0.050 0.009 0.274 4.538 0.092
4:1 0.089 0.080 0.009 0.063 0.042 0.021 0.257 4.495 0.082
8:1 0.125 0.116 0.009 0.084 0.045 0.039 0.323 5.285 0.115
Table 1: Ablation on RL-to-IL step mixing ratios in the reinforced post-training stage.
3.6 Auxiliary Objective
RL usually faces the challenge of sparse rewards, which makes the convergence process unstable and slow. To speed up convergence, we introduce auxiliary objectives that provide dense guidance to the entire action distribution.
The auxiliary objectives are designed to penalize undesirable behaviors by incorporating specific reward sources, including dynamic collisions, static collisions, positional deviations, and heading deviations. These objectives are computed based on the actions $a_{t}^{x,\text{old}}$ and $a_{t}^{y,\text{old}}$ selected by the old AD policy $\pi_{\theta_{\text{old}}}$ at time step $t$. To facilitate the evaluation of these actions, we separate the probability distribution of the action into four parts:
$f_{\text{dc}}=\begin{cases}1&\text{if the collision is ahead},\\ -1&\text{if the collision is behind}.\end{cases}$ (10)
$\mathcal{L}_{\text{dc}}(\theta_{y})=\mathbb{E}_{t}\left[\hat{A}_{t}^{\text{dc}}\cdot f_{\text{dc}}\cdot(\Delta\pi_{y}^{\text{dec}}-\Delta\pi_{y}^{\text{acc}})\right],$ (11)
$f_{\text{sc}}=\begin{cases}1&\text{if static obstacle is on the left},\\ -1&\text{if static obstacle is on the right}.\end{cases}$ (12)
$\mathcal{L}_{\text{sc}}(\theta_{x})=\mathbb{E}_{t}\left[\hat{A}_{t}^{\text{sc}}\cdot f_{\text{sc}}\cdot(\Delta\pi_{x}^{\text{right}}-\Delta\pi_{x}^{\text{left}})\right],$ (13)
$f_{\text{pd}}=\begin{cases}1&\text{if ego vehicle deviates leftward},\\ -1&\text{if ego vehicle deviates rightward}.\end{cases}$ (14)
The auxiliary objective for positional deviation correction is:
$\mathcal{L}_{\text{pd}}(\theta_{x})=\mathbb{E}_{t}\left[\hat{A}_{t}^{\text{pd}}\cdot f_{\text{pd}}\cdot(\Delta\pi_{x}^{\text{right}}-\Delta\pi_{x}^{\text{left}})\right],$ (15)
$f_{\text{hd}}=\begin{cases}1&\text{if ego vehicle deviates clockwise},\\ -1&\text{if ego vehicle deviates counterclockwise}.\end{cases}$ (16)
$\mathcal{L}_{\text{hd}}(\theta_{x})=\mathbb{E}_{t}\left[\hat{A}_{t}^{\text{hd}}\cdot f_{\text{hd}}\cdot(\Delta\pi_{x}^{\text{right}}-\Delta\pi_{x}^{\text{left}})\right],$ (17)
[…]
Collision Collision Deviation Deviation
@@ 3.6 Auxiliary Objective
Table 2: Ablation on reward sources. The table shows the impact of different reward components on performance.
@@
Auxiliary Obj. Auxiliary Obj. Auxiliary Obj. Auxiliary Obj.
@@
Table 3: Ablation on auxiliary objectives. The table shows the impact of different auxiliary objectives on performance.
Overall Auxiliary Objectives. The overall auxiliary objectives are a weighted sum of the individual objectives:
$\mathcal{L}_{\text{aux}}(\theta)=$ $\lambda_{1}\mathcal{L}_{\text{dc}}(\theta_{y})+\lambda_{2}\mathcal{L}_{\text{sc}}(\theta_{x})+$ (18)
$\lambda_{3}\mathcal{L}_{\text{pd}}(\theta_{x})+\lambda_{4}\mathcal{L}_{\text{hd}}(\theta_{x}),$
Optimization Objective. The final optimization objective combines the clipped PPO objective with the auxiliary objective:
$\mathcal{L}(\theta)=\mathcal{L}^{\text{PPO}}(\theta)+\mathcal{L}_{\text{aux}}(\theta).$ (19)
@@ 4.1 Experimental Settings
Dataset and Benchmark. We collect $2000h$ of expert human driving demonstrations in the real physical world. We get ground-truths of maps and agents in these driving demonstrations through a low-cost automated annotation pipeline. We use the map and agent labels as supervision for the first-stage perception pre-training. For the second-stage planning pre-training, we use the odometry information of the ego vehicle as supervision. For the third-stage reinforced post-training, we select 4305 critical dense-traffic clips of high collision risks from collected driving demonstrations and reconstruct these clips into 3DGS environments. Of these, 3968 3DGS environments are used for RL training, and the other 337 3DGS environments are used as closed-loop evaluation benchmarks.
Metric. We evaluate the performance of the AD policy using nine key metrics. Dynamic Collision Ratio (DCR) and Static Collision Ratio (SCR) quantify the frequency of collisions with dynamic and static obstacles, respectively, with their sum represented as the Collision Ratio (CR). Positional Deviation Ratio (PDR) measures the ego vehicle’s deviation from the expert trajectory with respect to position, while Heading Deviation Ratio (HDR) evaluates the ego vehicle’s consistency to the expert trajectory with respect to the forward direction. The overall deviation is quantified by the Deviation Ratio (DR), defined as the sum of PDR and HDR. Average Deviation Distance (ADD) quantifies the mean closest distance between the ego vehicle and the expert trajectory before any collisions or deviations occur. Additionally, Longitudinal Jerk (Long. Jerk) and Lateral Jerk (Lat. Jerk) assess driving smoothness by measuring acceleration changes in the longitudinal and lateral directions. CR, DCR, and SCR mainly reflect the policy’s safety, and ADD reflects the trajectory consistency between the AD policy and human drivers.
@@
VAD [17] 0.335 0.273 0.062 0.314 0.255 0.059 0.304 5.284 0.550
GenAD [49] 0.341 0.299 0.042 0.291 0.160 0.131 0.265 11.37 0.320
VADv2 [2] 0.270 0.240 0.030 0.243 0.139 0.104 0.273 7.782 0.171
@@
Table 4: Closed-loop quantitative comparisons with other IL-based methods on the 3DGS dense-traffic evaluation benchmark.
Figure 5: Closed-loop qualitative comparisons between the IL-only policy and RAD in 3DGS environments. Rows 1-2 correspond to Yield to Pedestrians. Rows 3-4 correspond to Unprotected Left-turn.
@@ 4.2 Ablation Study
To evaluate the impact of different design choices in RAD, we conduct three ablation studies. These studies examine the balance between RL and IL, the role of different reward sources, and the effect of auxiliary objectives.
RL-IL Ratio Analysis. We first analyze the effect of different RL-to-IL step mixing ratios (Tab. 1). A pure IL policy (0:1) results in the highest CR (0.229) but the lowest ADD (0.238), indicating strong trajectory consistency but poor safety. In contrast, a pure RL policy (1:0) significantly reduces CR (0.143) but increases ADD (0.345), suggesting improved safety at the cost of trajectory deviation. The best balance is achieved at a 4:1 ratio, which results in the lowest CR (0.089) while maintaining a relatively low ADD (0.257). Further increasing RL dominance (e.g., 8:1) leads to a deteriorated ADD (0.323) and higher jerk, implying reduced trajectory smoothness.
@@
Auxiliary Objective Analysis. Finally, we examine the impact of auxiliary objectives (Tab. 3). Compared to the full auxiliary setup (ID 8), omitting any auxiliary objective increases CR, with a significant rise observed when all auxiliary objectives are removed. This highlights their collective role in enhancing safety. Notably, ID 1, which retains all auxiliary objectives but excludes the PPO objective, results in a CR of 0.187. This value is higher than that of ID 8, indicating that while auxiliary objectives help reduce collisions, they are most effective when combined with the PPO objective.
Our ablation studies highlight the importance of combining RL and IL, using a comprehensive reward function, and implementing structured auxiliary objectives. The optimal RL-IL ratio (4:1) and the full reward and auxiliary setups consistently yield the lowest CR while maintaining stable ADD, ensuring both safety and trajectory consistency.
@@ 4.3 Comparisons with Existing Methods
As presented in Tab. 4, we compare RAD with other end-to-end autonomous driving methods in the proposed 3DGS-based closed-loop evaluation. For fair comparisons, all the methods are trained with the same amount of human driving demonstrations. The 3DGS environments for the RL training in RAD are also based on these data. RAD achieves better performance compared to IL-based methods in most metrics. Especially in terms of CR, RAD achieves 3×3$\times$3 × lower collision rate, demonstrating that RL helps the AD policy learn general collision avoidance ability.
@@ 4.4 Qualitative Comparisons
We provide qualitative comparisons between the IL-only AD policy (without reinforced post-training) and RAD, as shown in Fig. 5. The IL-only method struggles in dynamic environments, frequently failing to avoid collisions with moving obstacles or manage complex traffic situations. In contrast, RAD consistently performs well, effectively avoiding dynamic obstacles and handling challenging tasks. These results highlight the benefits of closed-loop training in the hybrid method, which enables better handling of dynamic environments. Additional visualizations are included in the Appendix (Fig. A1).
5 Limitation and Conclusion
In this work, we propose the first 3DGS-based RL framework for training end-to-end AD policy. We combine RL and IL, with RL complementing IL to model the causations and narrow the open-loop gap, and IL complementing RL in terms of human alignment. This work also has some limitations. Currently the used 3DGS environments are running in a non-reactive manner, i.e., other traffic participants do not react according to ego vehicle’s behavior but act with log replay. And the effect of 3DGS still has room for improvement, especially for rendering non-rigid pedestrians, unobserved views, and low-light scenarios. Future works will focus on solving these problems and scaling up RL to the next level.
Acknowledgement
We would like to acknowledge Qingjie Wang, Yongjun Yu, Zehua Li, Peng Wang, Nuoya Zhou, Songlin Yang, Ruiqi Wang, Tianheng Cheng, Changze Li, Zhe Chen, and Tong Qin for discussion and assistance.
Appendix A Appendix
A.1 Action Space Details
@@
$a^{x}\in\{d^{x}_{\text{min}},\dots,0,\dots,d^{x}_{\text{max}}\}.$ (A1)
@@
A.2 Implementation Details
@@
Planning Pre-Training. The action space is discretized using predefined anchors $\mathcal{A}=\{(a^{x}_{i},a^{y}_{j})\}_{i=1,j=1}^{N_{x},N_{y}}$. Each anchor corresponds to a specific steering-speed combination within the 0.5-second planning horizon. Given the ground truth vehicle position at $t=0.5\ \text{s}$ denoted as $p_{\text{gt}}=(p_{\text{gt}}^{x},p_{\text{gt}}^{y})$, we implement normalized nearest-neighbor matching over predefined anchor positions:
$\hat{i}$ $=\operatorname*{arg\,min}_{i}\left\|\frac{\mathbf{a}^{x}_{i}-d^{x}_{\text{min}}}{d^{x}_{\text{max}}-d^{x}_{\text{min}}}-\frac{p_{\text{gt}}^{x}-d^{x}_{\text{min}}}{d^{x}_{\text{max}}-d^{x}_{\text{min}}}\right\|_{2},$ (A3)
@@
Based on the matched anchor indices $(\hat{i},\hat{j})$, we formulate the imitation learning objective as a dual focal loss [29]:
$\mathcal{L}_{\text{IL}}=\mathcal{L}_{\text{focal}}(\pi(a^{x}\mid s),\hat{i}_{t})+\mathcal{L}_{\text{focal}}(\pi(a^{y}\mid s),\hat{j}_{t}),$ (A4)
where $\mathcal{L}_{\text{focal}}$ is focal loss for discrete action classification, and $\pi(a^{x}\mid s)$ and $\pi(a^{y}\mid s)$ are predicted action distributions from Eq. 1.
Training configurations. We provide detailed hyperparameters for the two main stages, Planning Pre-Training and Reinforced Post-Training, in Tab. A1 and Tab. A2, respectively.
@@
Table A1: Hyperparameters used in RAD Planning Pre-Training stage.
@@
Table A2: Hyperparameters used in RAD Reinforced Post-Training stage.
A.3 Metric Details
We evaluate the performance of the autonomous driving policy using eight key metrics.

### v2

@@ 3.6 Auxiliary Objective
Col. Col. Dev. Dev. $\downarrow Jerk\downarrow$
@@
Table 3: Impact of different auxiliary objectives on performance.
@@
Obj. Aux. Obj. Aux. Obj. Aux. Obj. Aux. Obj. $\downarrow Jerk\downarrow$
@@ 4.1 Experimental Settings
Dataset and Benchmark. We collect 2000 hours of expert driving demonstrations in real-world conditions and generate map and agent annotations via an automated pipeline for perception pre-training. Ego-vehicle odometry is employed for planning pre-training. For reinforcement learning, we select 4305 real-world driving scenes covering diverse road types, traffic densities, and agent behaviors to ensure environmental diversity. These scenes are first reconstructed into 3DGS environments, from each of which a fixed-length 8-second clip is extracted. Among these clips, 3968 are used for RL training, and the other 337 are used as closed-loop evaluation benchmarks. More training details are in the Appendix A.5.
Table 4: Closed-loop quantitative comparisons with other IL-based methods on the 3DGS evaluation benchmark.
@@
$\downarrow Jerk\downarrow$
TransFuser transfuser 0.320 0.273 0.047 0.235 0.188 0.047 0.263 4.538 0.142
VAD vad 0.335 0.273 0.062 0.314 0.255 0.059 0.304 5.284 0.550
GenAD zheng2024genad 0.341 0.299 0.042 0.291 0.160 0.131 0.265 11.37 0.320
VADv2 vadv2 0.270 0.240 0.030 0.243 0.139 0.104 0.273 7.782 0.171
@@
Metric. We evaluate the performance of the AD policy using nine key metrics. Dynamic Collision Ratio (DCR) and Static Collision Ratio (SCR) quantify the frequency of collisions with dynamic and static obstacles, respectively, with their sum represented as the Collision Ratio (CR). Positional Deviation Ratio (PDR) measures the ego vehicle’s deviation from the expert trajectory with respect to position, while Heading Deviation Ratio (HDR) evaluates the ego vehicle’s consistency to the expert trajectory with respect to the forward direction. The overall deviation is quantified by the Deviation Ratio (DR), defined as the sum of PDR and HDR. Average Deviation Distance (ADD) quantifies the mean closest distance between the ego vehicle and the expert trajectory before any collisions or deviations occur. Additionally, Longitudinal Jerk (Long. Jerk) and Lateral Jerk (Lat. Jerk) assess driving smoothness by measuring acceleration changes in the longitudinal and lateral directions. CR, DCR, and SCR mainly reflect the policy’s safety, and ADD reflects the trajectory consistency between the AD policy and human drivers. More details are provided in Appendix A.6.
@@ 4.2 Ablation Study
To evaluate the impact of different design choices in RAD, we conduct three ablation studies. These studies highlight the importance of combining RL and IL, the role of different reward sources, and the effect of auxiliary objectives.
Training Strategy Comparison. We compare pure IL, pure RL, and our mixed IL+RL strategy in reinforced post-training (Tab. 1). Pure IL achieves low deviation (ADD 0.238) but suffers from high collision risk (CR 0.229). Pure RL improves safety (CR 0.143) but deviates more from expert behavior (ADD 0.345). Our IL+RL strategy achieves the best balance (CR 0.089, ADD 0.257), demonstrating improved safety without sacrificing behavioral fidelity.
@@
Auxiliary Objective Analysis. We examine the impact of auxiliary objectives (Tab. 3). Compared to the full auxiliary setup (ID 8), omitting any auxiliary objective increases CR, with a significant rise observed when all auxiliary objectives are removed. This highlights their collective role in enhancing safety. Notably, ID 1, which retains all auxiliary objectives but excludes the PPO objective, results in a CR of 0.187. This value is higher than that of ID 8, indicating that while auxiliary objectives help reduce collisions, they are most effective when combined with the PPO objective.
@@ 4.3 Comparisons with Existing Methods
As presented in Tab. 4, we compare RAD with other end-to-end autonomous driving methods in the proposed 3DGS-based closed-loop evaluation. For fair comparisons, all the methods are trained with the same amount of human driving demonstrations. The 3DGS environments for the RL training in RAD are also based on these data. RAD achieves better performance compared to IL-based methods in most metrics. Especially in terms of CR, RAD achieves $3\times$ lower collision rate, demonstrating that RL helps the AD policy learn general collision avoidance ability.
Figure 5: Consistency analysis between 3DGS environment and real-world environment. We compare the driving behaviors of the same driving policy in both environments. Subfigure (1)–(4) present the first and last frames during closed-loop evaluation in both environments. Subfigure (5)-(6) depict the temporal evolution of the ego vehicle’s position.
4.4 Consistency Analysis
To demonstrate the consistency between 3DGS environment and real-world environment, we compare the driving behaviors of the same driving policy in both environments. As illustrated in Fig. 5, subfigure (1)–(4) present the first and last frames during closed-loop evaluation in both environments. Subfigure (5)-(6) depict the temporal evolution of the ego vehicle’s position. Both qualitative and quantitative results indicate a high degree of behavioral consistency between the 3DGS environment and the real world environment. It shows 3DGS-based closed-loop evaluation can reflect the real-world driving performance.
Figure 6: Closed-loop qualitative comparisons between the IL-only policy and RAD in a Yield to Pedestrians scenario. The IL-only policy fails to yield (Row 1), while RAD successfully yields to pedestrians (Row 2).
@@ 4.5 Qualitative Comparisons
We provide qualitative comparisons between the IL-only AD policy (without reinforced post-training) and RAD, as shown in Fig. 6. The IL-only method struggles in dynamic environments, frequently failing to avoid collisions with moving obstacles or manage complex traffic situations. In contrast, RAD consistently performs well, effectively avoiding dynamic obstacles and handling challenging tasks. These results highlight the benefits of closed-loop training in the hybrid method, which enables better handling of dynamic environments. Additional visualizations are included in Fig. 7.
5 Conclusion
In this work, we propose RAD, the first 3DGS-based closed-loop reinforcement learning framework for end-to-end autonomous driving. We combine RL and IL, with RL complementing IL to model the causations and narrow the open-loop gap, and IL complementing RL in terms of human alignment. Complemented by a targeted reward system and auxiliary objectives, RAD achieves 3× lower collision rates than state-of-the-art IL methods, with strong performance in challenging scenarios like unprotected left-turns and dense traffic.
Limitations and future work. The effect of 3DGS still has room for improvement, particularly in rendering non-rigid pedestrians, unobserved views, and low-light scenarios. Future works will focus on addressing these issues and scaling up RL to the next level.
Acknowledgement: This work was partially supported by National Natural Science Foundation of China (No. 62376102).
Appendix A Technical Appendices
A.1 3DGS Reconstruction and Rendering Optimization
To support closed-loop training, we extend the StreetGaussian [20] framework, focusing on enhancing rendering realism and geometric accuracy, particularly for high-fidelity rendering in off-trajectory views.
First, we employ mesh modeling to constrain the geometry of the road surface, constraining Gaussian spheres to the mesh surface to ensure precise road geometry from any viewpoint. Additionally, we model the sky separately to prevent confusion with foreground objects, improving rendering realism under complex lighting conditions.
For foreground objects (e.g., vehicles, pedestrians), we optimize their poses during training and incorporate depth and normal consistency as supervision signals to further enhance geometric reconstruction accuracy and surface detail fidelity. These optimizations significantly improve rendering quality in novel viewpoints, especially in dynamic scenes, where object motion trajectories and surface details remain consistent with real-world observations.
These improvements allow the 3DGS environment to more effectively support closed-loop training, providing a foundation of high realism and geometric accuracy for large-scale trial-and-error learning in autonomous driving strategies.
A.2 Action Space Details
@@
$a^{x}\in\{d^{x}_{\text{min}},\dots,0,\dots,d^{x}_{\text{max}}\}.$ (9)
@@
A.3 Policy Optimization Details
GAE Computation Details. The advantage estimates $\hat{A}_{t}^{x}$ and $\hat{A}_{t}^{y}$ are then computed as follows:
$\delta_{t}^{x}$ $=r_{t}^{x}+\gamma V_{x}(s_{t+1})-V_{x}(s_{t}),$ (11)
$\delta_{t}^{y}$ $=r_{t}^{y}+\gamma V_{y}(s_{t+1})-V_{y}(s_{t}),$
$\hat{A}_{t}^{x}$ $=\sum_{l=0}^{\infty}(\gamma\lambda)^{l}\delta_{t+l}^{x},$
$\hat{A}_{t}^{y}$ $=\sum_{l=0}^{\infty}(\gamma\lambda)^{l}\delta_{t+l}^{y},$
Complete PPO Objectives. The full clipping objectives are:
A.4 Auxiliary Objectives Details
Action Probability Decomposition. The auxiliary objectives are designed to penalize undesirable behaviors by incorporating specific reward sources, including dynamic collisions, static collisions, positional deviations, and heading deviations. These objectives are computed based on the actions $a_{t}^{x,\text{old}}$ and $a_{t}^{y,\text{old}}$ selected by the old AD policy $\pi_{\theta_{\text{old}}}$ at time step $t$. To facilitate the evaluation of these actions, we separate the probability distribution of the action into four parts:
[…]
$\mathcal{L}_{\text{dc}}(\theta)=\mathbb{E}_{t}\left[\hat{A}_{t}^{\text{dc}}\cdot f_{\text{dc}}\cdot(\Delta\pi_{y}^{\text{dec}}-\Delta\pi_{y}^{\text{acc}})\right],$ (15)
[…]
$\mathcal{L}_{\text{sc}}(\theta)=\mathbb{E}_{t}\left[\hat{A}_{t}^{\text{sc}}\cdot f_{\text{sc}}\cdot(\Delta\pi_{x}^{\text{right}}-\Delta\pi_{x}^{\text{left}})\right],$ (17)
[…]
The auxiliary objective for positional deviation correction is:
$\mathcal{L}_{\text{pd}}(\theta)=\mathbb{E}_{t}\left[\hat{A}_{t}^{\text{pd}}\cdot f_{\text{pd}}\cdot(\Delta\pi_{x}^{\text{right}}-\Delta\pi_{x}^{\text{left}})\right],$ (19)
[…]
$\mathcal{L}_{\text{hd}}(\theta)=\mathbb{E}_{t}\left[\hat{A}_{t}^{\text{hd}}\cdot f_{\text{hd}}\cdot(\Delta\pi_{x}^{\text{right}}-\Delta\pi_{x}^{\text{left}})\right],$ (21)
A.5 Implementation Details
@@
Planning Pre-Training. The action space is discretized using predefined anchors $\mathcal{A}=\{(a^{x}_{i},a^{y}_{j})\}_{i=1,j=1}^{N_{x},N_{y}}$. Each anchor corresponds to a specific steering-speed combination within the 0.5-second planning horizon. Given the ground truth vehicle position at t=0.5st=0.5\ $\text${s} denoted as $p_{\text{gt}}=(p_{\text{gt}}^{x},p_{\text{gt}}^{y})$, we implement normalized nearest-neighbor matching over predefined anchor positions:
$\hat{i}$ $=\operatorname*{arg\,min}_{i}\left\|\frac{\mathbf{a}^{x}_{i}-d^{x}_{\text{min}}}{d^{x}_{\text{max}}-d^{x}_{\text{min}}}-\frac{p_{\text{gt}}^{x}-d^{x}_{\text{min}}}{d^{x}_{\text{max}}-d^{x}_{\text{min}}}\right\|_{2},$ (22)
@@
[…]
$\mathcal{L}_{\text{IL}}=\mathcal{L}_{\text{focal}}(\pi(a^{x}\mid s),\hat{i}_{t})+\mathcal{L}_{\text{focal}}(\pi(a^{y}\mid s),\hat{j}_{t}),$ (23)
where $\mathcal{L}_{\text{focal}}$ is focal loss for discrete action classification.
Reinforced Post-Training. During the training process, we use a cycle where reinforcement learning (RL) and imitation learning (IL) alternate. In each full cycle, we run four rounds of RL training, followed by one round of IL training.
Each RL training round consists of 320 iterations. Our clips are 8 seconds long with a frame rate of 10Hz, meaning that 320 iterations cover four clips of data. We employ a sliding window mechanism that holds four clips of data, updated in a first-in, first-out manner. In each RL training round, data from one new clip is collected and the sliding window is updated accordingly. We find that an RL-to-IL ratio of $(320\times 4):320=4:1$ yields the best results.
Training configurations. We provide detailed hyperparameters for the two main stages, Planning Pre-Training and Reinforced Post-Training, in Tab. 5 and Tab. 6, respectively.
Table 5: Hyperparameters used in RAD Planning Pre-Training stage.
@@
Traning GPU 128 RTX4090
Table 6: Hyperparameters used in RAD Reinforced Post-Training stage.
@@
Traning GPU 32 RTX4090
A.6 Metric Details
We evaluate the performance of the autonomous driving policy using nine key metrics.